# Bridging Synthetic and Real Routing Problems via LLM-Guided Instance Generation and Progressive Adaptation

### v1

@@ Appendix B Evolution Mechanics
Evolution Mechanisms for Creating New Generators. We build our evolution framework upon the structure of ReEvo (ye2024reevo), with modifications on several key components for better evolution-search performance. Evolution mechanisms can be classified into two groups: Searching (crossover, mutation) and Reflection (short-term reflection, long-term reflection), and they are arranged into the following sequential order in one evolution iteration:
@@ Appendix C Routing Problem Formulation
CVRP. (10.5555/2723809) Every node in CVRP has two elements: node coordinate $v_{i}$ and demand $\delta_{i}$ for $n\in\{1,2,3,...n\}$. Specifically, the demand $\delta_{0}$ for the depot node $v_{0}$ is 0. The solution to the problem consists of multiple subtours, each starting and terminating at the depot. A feasible tour should satisfy: 1) every customer node in the problem is visited only once. 2) The total demand in each subtour does not exceed the pre-defined vehicle capacity.
In this paper, we mainly consider the evolution of the distributions of the customer nodes, which is consistent with the design of our framework. To craft the seed generator for CVRP data evolution, we also categorize CVRP instances into three sets according to different depot positioning schemes (GOUVEIA1996178): 1) S1: depot is located in the center of the unit square. 2) S2: depot is randomly sampled from the unit square $0,1]^{2}$. 3) S3: depot is at (0,0). Nonetheless, instead of evolving the generator of each category one at a time, we incorporate these three types of depots into one generator to save evolution cost. Following existing works (uchoa2017new; NEURIPS2018_9fb4651c), the coordinates of the customer nodes are sampled uniformly from the unit square while the demand $d_{i}$ of each customer node is initialized with $U[1,10]$. The capacity $Q$ of each vehicle is set according to the following formula: $\mathrm{Q}=\max\left(\left\lceil r\cdot\frac{1}{n}\sum_{i=1}^{n}d_{i}\right\rceil,\,\left\lceil k\cdot\max_{1\leq i\leq n}d_{i}\right\rceil\right)$ where $r\sim\mathrm{Triangular}(3,\,6,\,25)$ and $k$ is a constant (e.g. $k=2$). The demands and capacity are further normalized to $\delta_{i}^{\prime}=\delta_{i}/Q$ and 1, respectively.
@@ Appendix D Experimental Setup
Our framework utilizes the identical network topologies of POMO (NEURIPS2020_f231f210) and LEHD (luo2023neural). For proxy evaluation in LLM-guided evolution and phase one of progressive fine-tuning, we employed the pre-trained TSP and CVRP models trained with problems of size 100, using their released checkpoints. For proxy evaluation and phase one fine-tuning, we also generate TSP100 and CVRP100 for training. Additionally, we employ the same data loader as POMO’s and LEHD’s. Specifically, for fine-tuning LEHD in phase one, training instances are generated only once and shuffled for reuse when training every epoch. For the results in Table 1 and Table 2, we set the hyperparameters in Table 7 for LLM-guided evolution and proxy evaluation.
@@
Proxy-evaluation & Phase One Fine-tuning Setup. For proxy-evaluation in LLM-based evolution and phase one fine-tuning, we maintain most of the hyperparameters the same as the original experimental settings in the respective model papers: Adam (kingma2014adam) optimizer is used with an initial learning rate of $10^{-4}$ and a decay of 0.9 for both TSP and CVRP, except for the LEHD TSP model, where a lower learning rate of $10^{-5}$ and a decay of 0.6 is used. We use the same batch sizes as the original papers. However, for phase one fine-tuning, we fine-tuned the pre-trained models with more epochs and instances to see the convergence. For POMO in phase one, the model for TSP is fine-tuned for 300 epochs with 5000 problems in each episode, while the model for CVRP is fine-tuned for 200 epochs with 500 problems in each episode. For LEHD in phase one, we fine-tuned their pre-trained TSP version and their CVRP version both for 20 epochs. And in each episode for LEHD, we generate 50000 synthetic data once for the repetitive usage in each epoch. For finetuning both models in phase one, we generate mixed distributions of S1, S2, S3 according to the ratio of S1, S2, S3 problems in the validation set in one epoch (e.g. if S1 problems: S2 problems: S3 problems= 17:19:12, then the ratio of the synthetic instances of type S1, S2, S3 is also 17:19:12).
@@ Appendix E Additional Results
To further evaluate generalization ability of the models trained with our framework, particularly on the data unseen during the generator evolution and progressive fine-tuning, we tested them directly on 22 held-out problems. The 22 unseen problems include 5 small-size instances and 17 instances of sizes n $\in$ $[1300,5000)]$ proposed in TSPLib95 (reinelt1991tsplib). Original POMO (NEURIPS2020_f231f210) and LEHD (luo2023neural), along with those trained from progressive fine-tuning, are employed to solve the instances, with both models tested under their augmentation version (×8 augmentation or RRC-50). We report the optimum, the objective values, and the gaps for each problem in Table 6. As shown, LEHD trained from progressive finetuning (i.e., LEHD (ours) in the table) outperforms other models in most problems, especially on large-size ones like pcb3038 and fnl4461, while it is still comparable to the original model on the first five small instances. As for POMO (ours), all problems have seen reductions in gaps by $0.92\%~\sim~60\%+$. These results demonstrate that the models with our methods, despite not being exposed to these unseen TSPLib instances during fine-tuning, consistently achieve strong generalization performance, particularly on large-scale and challenging cases. This improvement can be largely attributed to the structural prior alignment introduced in the first phase of progressive fine-tuning, which enables the model to internalize diverse distributional patterns and transfer this knowledge to previously unseen problem instances.
@@ Appendix F Prompts
Prompts adopted for EvoReal are gathered in this section. For evolution operators (crossover and mutation) and reflection operators (long-term reflection and short-term reflection), since we only made slight modifications to those of ReEvo (ye2024reevo), these prompts are not listed here.

### v2

@@ Appendix B Evolution Mechanics
Evolution Mechanisms for Creating New Generators. We build our evolution framework upon the structure of ReEvo (Ye et al. 2024), with modifications on several key components for better evolution-search performance. Evolution mechanisms can be classified into two groups: Searching (crossover, mutation) and Reflection (short-term reflection, long-term reflection), and they are arranged into the following sequential order in one evolution iteration:
@@ Appendix C Routing Problem Formulation
CVRP. (Toth et al. 2014) Every node in CVRP has two elements: node coordinate $v_{i}$ and demand $\delta_{i}$ for $n\in\{1,2,3,...n\}$. Specifically, the demand $\delta_{0}$ for the depot node $v_{0}$ is 0. The solution to the problem consists of multiple subtours, each starting and terminating at the depot. A feasible tour should satisfy: 1) every customer node in the problem is visited only once. 2) The total demand in each subtour does not exceed the pre-defined vehicle capacity.
In this paper, we mainly consider the evolution of the distributions of the customer nodes, which is consistent with the design of our framework. To craft the seed generator for CVRP data evolution, we also categorize CVRP instances into three sets according to different depot positioning schemes (Gouveia 1996): 1) S1: depot is located in the center of the unit square. 2) S2: depot is randomly sampled from the unit square $0,1]^{2}$. 3) S3: depot is at (0,0). Nonetheless, instead of evolving the generator of each category one at a time, we incorporate these three types of depots into one generator to save evolution cost. Following existing works (Uchoa et al. 2017; Nazari et al. 2018), the coordinates of the customer nodes are sampled uniformly from the unit square while the demand $d_{i}$ of each customer node is initialized with $U[1,10]$. The capacity $Q$ of each vehicle is set according to the following formula: $\mathrm{Q}=\max\left(\left\lceil r\cdot\frac{1}{n}\sum_{i=1}^{n}d_{i}\right\rceil,\,\left\lceil k\cdot\max_{1\leq i\leq n}d_{i}\right\rceil\right)$ where $r\sim\mathrm{Triangular}(3,\,6,\,25)$ and $k$ is a constant (e.g. $k=2$). The demands and capacity are further normalized to $\delta_{i}^{\prime}=\delta_{i}/Q$ and 1, respectively.
@@ Appendix D Experimental Setup
Our framework utilizes the identical network topologies of POMO (Kwon et al. 2020) and LEHD (Luo et al. 2023). For proxy evaluation in LLM-guided evolution and phase one of progressive fine-tuning, we employed the pre-trained TSP and CVRP models trained with problems of size 100, using their released checkpoints. For proxy evaluation and phase one fine-tuning, we also generate TSP100 and CVRP100 for training. Additionally, we employ the same data loader as POMO’s and LEHD’s. Specifically, for fine-tuning LEHD in phase one, training instances are generated only once and shuffled for reuse when training every epoch. For the results in Table 1 and Table 2, we set the hyperparameters in Table 7 for LLM-guided evolution and proxy evaluation.
@@
Proxy-evaluation & Phase One Fine-tuning Setup. For proxy-evaluation in LLM-based evolution and phase one fine-tuning, we maintain most of the hyperparameters the same as the original experimental settings in the respective model papers: Adam (Kingma and Ba 2014) optimizer is used with an initial learning rate of $10^{-4}$ and a decay of 0.9 for both TSP and CVRP, except for the LEHD TSP model, where a lower learning rate of $10^{-5}$ and a decay of 0.6 is used. We use the same batch sizes as the original papers. However, for phase one fine-tuning, we fine-tuned the pre-trained models with more epochs and instances to see the convergence. For POMO in phase one, the model for TSP is fine-tuned for 300 epochs with 5000 problems in each episode, while the model for CVRP is fine-tuned for 200 epochs with 500 problems in each episode. For LEHD in phase one, we fine-tuned their pre-trained TSP version and their CVRP version both for 20 epochs. And in each episode for LEHD, we generate 50000 synthetic data once for the repetitive usage in each epoch. For finetuning both models in phase one, we generate mixed distributions of S1, S2, S3 according to the ratio of S1, S2, S3 problems in the validation set in one epoch (e.g. if S1 problems: S2 problems: S3 problems= 17:19:12, then the ratio of the synthetic instances of type S1, S2, S3 is also 17:19:12).
@@ Appendix E Additional Results
To further evaluate generalization ability of the models trained with our framework, particularly on the data unseen during the generator evolution and progressive fine-tuning, we tested them directly on 22 held-out problems. The 22 unseen problems include 5 small-size instances and 17 instances of sizes n $\in$ $[1300,5000)]$ proposed in TSPLib95 (Reinelt 1991). Original POMO (Kwon et al. 2020) and LEHD (Luo et al. 2023), along with those trained from progressive fine-tuning, are employed to solve the instances, with both models tested under their augmentation version (×8 augmentation or RRC-50). We report the optimum, the objective values, and the gaps for each problem in Table 6. As shown, LEHD trained from progressive finetuning (i.e., LEHD (ours) in the table) outperforms other models in most problems, especially on large-size ones like pcb3038 and fnl4461, while it is still comparable to the original model on the first five small instances. As for POMO (ours), all problems have seen reductions in gaps by $0.92\%~\sim~60\%+$. These results demonstrate that the models with our methods, despite not being exposed to these unseen TSPLib instances during fine-tuning, consistently achieve strong generalization performance, particularly on large-scale and challenging cases. This improvement can be largely attributed to the structural prior alignment introduced in the first phase of progressive fine-tuning, which enables the model to internalize diverse distributional patterns and transfer this knowledge to previously unseen problem instances.
@@ Appendix F Prompts
Prompts adopted for EvoReal are gathered in this section. For evolution operators (crossover and mutation) and reflection operators (long-term reflection and short-term reflection), since we only made slight modifications to those of ReEvo (Ye et al. 2024), these prompts are not listed here.